OpenStudy (mathmath333):

Due to a 25 % increase in the price of rice per kilogram a person is able to purchase 20 kilograms less for Rs 400. What is the increased price of rice per kilogram? and also give orignal price

OpenStudy (perl):

what work do you have so far?

OpenStudy (perl):

you are able to purchase 20 kilograms less for 400 Rs, not sure what R stands for

OpenStudy (mathmath333):

m unable to form the equation

OpenStudy (perl):

Rupee?

OpenStudy (mathmath333):

Rs--> currency

OpenStudy (perl):

you can make a proportion

OpenStudy (mathmath333):

how

OpenStudy (perl):

the draw button does not work

OpenStudy (mathmath333):

u can tell oraly

OpenStudy (mathmath333):

@ikram002p

OpenStudy (perl):

try this, Let x = original price 1.25x = 400 x = 400 / 1.25

OpenStudy (perl):

original price x = 320 and the increase in price is 80 Rs

OpenStudy (mathmath333):

but this not price per kilo gram? right

OpenStudy (perl):

this is the price for the original amount (more 20 kg )

OpenStudy (mathmath333):

ok but how will u get price/kg

OpenStudy (perl):

there is a difference of 80 dollars, between the old price and the new price

OpenStudy (perl):

and a difference of 20 kilograms

OpenStudy (perl):

so 80 Rs / 20 kg = 4 Rs / kg price increase

OpenStudy (mathmath333):

this is orignal price?

OpenStudy (perl):

no, the english is very important here

OpenStudy (perl):

What is the increased price of rice per kilogram? 4 Rs/ kg and also give original price: 320 Rs

OpenStudy (perl):

we don't have to find the original price of rice per kilogram

OpenStudy (perl):

i don't think we can find that, since we are not given the original amount of kg of rice

OpenStudy (perl):

All you are given is that Due to a 25 % increase in the price of rice per kilogram you are able to purchase 20 kg less than what you usually buy for Rs 400. The number of kg you usually buy is not given

OpenStudy (perl):

ok i think we are done here, let me know if you have any questions

OpenStudy (mathmath333):

but the majority gave answer as Rs 5/-

OpenStudy (mathmath333):

and the initial price is given Rs 4/-

OpenStudy (mathmath333):

in book

OpenStudy (perl):

what was the book answer?

OpenStudy (perl):

the question does not ask for initial price. it asks what is the increase in price per kilogram of rice. the increase in price is 4 Rs

OpenStudy (perl):

1 Indian Rupee equals 0.016 US Dollar

OpenStudy (perl):

4 Indian Rupee equals 0.065 US Dollar so the price increase was 6.5 cents for a kilogram of rice

OpenStudy (perl):

6 pennies and half a penny :)

OpenStudy (mathmath333):

since the price is rising by 25%, consuption has to decrease by 20% but their is an actual reduction in consuption by 20 kg hence the orignal price is 100kg money spent being RS 400 ,the orignal price of rice is 4kg

OpenStudy (perl):

price is not kg, thats weight

OpenStudy (mathmath333):

*RS 4/kg

OpenStudy (perl):

why does consumption have to decrease by 20% ? i might be missing information about the problem

OpenStudy (perl):

I only answered based on what I was given

OpenStudy (perl):

sorry, what was the book answer exactly ?

OpenStudy (perl):

I thought this was an algebra question

OpenStudy (mathmath333):

i also dont know consuption factor here

OpenStudy (perl):

hmm

OpenStudy (perl):

a rise in 6 cents probably wont stop a person from purchasing the same amount of rice

OpenStudy (mathmath333):

book gave only orignal price , i have to find increased price

OpenStudy (perl):

here is the answer too http://totalgadha.com/mod/forum/discuss.php?d=8885

OpenStudy (mathmath333):

here is i found on somesite \("\)saving due to price reduction= 25% of 400 is 25* 400/100 this savings enable him to buy 20kg and, cost of 20kg at the reduced price= 25*400/100 then cost of 1kg at the reduced price= 25*400/100*20 = Rs 5\(" \) but i dont undertand this

OpenStudy (perl):

one moment, reading

OpenStudy (perl):

it seems like theres two different answers

OpenStudy (perl):

if you assume that consumption has to decrease, then you get a different answer

OpenStudy (mathmath333):

so which one is correct btw the link u gave gives initial price

OpenStudy (perl):

i guess it depends on the assumptions of the person who made the question. how did they figure out by what percentage consumption has to decrease?

OpenStudy (perl):

this website also gets 4 Rs http://mathbyvemuri.blogspot.com/2012/10/deatiled-solutions-for-competitive-exam.html

OpenStudy (mathmath333):

they assumed 20 kg to be 20% decrease

OpenStudy (perl):

so they assumed he initially bought 100 kg?

OpenStudy (mathmath333):

but this site gives RS 5/- http://www.pagalguy.com/discussions/cgl-2013-tier-2-preparation-25099757/17129618

OpenStudy (perl):

thats a book

OpenStudy (perl):

i cant click that, it says i have to log in.

OpenStudy (mathmath333):

ok let me give u answers frm that

OpenStudy (mathmath333):

Vibhor Sharma @ScareCrow2812.7k Price ratio = 1.25 = 5:4 Consumption ratio = 4:5 = 1/5 of original less than original 1/5 ~ 20 Kg 1 ~ 100 Kg 100 Kg ~ Rs 400 1 Kg ~ Rs 4 New Price = 4*1.25 = 5 Reply • Like •5 Like 04 Aug '13 •#1 @Harmeet89 Harmeet Matharu @Harmeet896.8k A rise of 25% in the cost corresponds to a fall of 20% in the consumption, since the expenditure on the purchase is the same. Hence he gets 20% less which corresponds to 20kg less rice than the original. Hence he gets 80 kgs for 400 Rs. Hence new price = 400/80 = 5 Rs. Reply • Like •3 Like 04 Aug '13 •#2 @Rajdeep_S Raj S @Rajdeep_S7.1k 25% of 400 is 100 100 Rs - 20 kg 5 rs - 1kg P.S. - Don't trust my methods. Reply • Like •5 Like 04 Aug '13 •#3 @vee_key jitender kumar @vee_key288 saving due to price reduction= 25% of 400 is 25* 400/100 this savings enable him to buy 20kg and, cost of 20kg at the reduced price= 25*400/100 then cost of 1kg at the reduced price= 25*400/100*20 = Rs 5 Reply • Like 04 Aug '13 •#4 @coolrakesh cgl2014 out @coolrakesh1.8k short cut---- If a man purchase X kg less in Y rs due to rate increased by R% then New Price of Product = (R/100) * Y/X,,, Old Price= (R / 100+R) * Y/X.... Reply • Like 04 Aug '13 •#5 @Anonymous @Anonymous price up by 1/4, consumption down by 1/5, 1/5 = 20, 1 = 100 kg, old price = 400/100 = 4 new price = 400/80 = 5 Reply • Like •6 Like 04 Aug '13 •#6 @manash9078 newton @manash90781.5k if increase in price is 1/4(i.e. 25%), decrease in consumption= 1/5 so, (1/5)x=20 x=100 so, original price is Rs. 400/100=Rs.4 so, new price is rs.(4+1)=5